OpenStudy (anonymous):

determine whether the ratios are propotional: 25, 30 -- -- 40 48

OpenStudy (anonymous):

they are both equal to 5/8.

OpenStudy (anonymous):

Yes they are. If you break them down the the lowest factors they both equal 5/8.

OpenStudy (anonymous):

can u explain how u got that?

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

25/40 Divide both by 5 = 5/8 30/48 Divide both by 6 = 5/8

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

c: